If nC2=nC3,calculate the value of 6Cn-1
Answer by stanbon(75887)

If nC2=nC3,calculate the value of 6Cn-1
-----
In general nCr = nC(n-r)
---------------------------------
If nC2 = nC3, 3 = n-2 and n = 5
-----
Then 6Cn-1 = 6C(5-1) = 6C4 = 6C2 = (6*5)/(1*2) = 15
-------------
Cheers,
Stan H.
-----------